OpenStudy (anonymous):

Find the slope and the y-intercept of the given line y=4x+11

OpenStudy (anonymous):

that equation is in the slop intercept form

OpenStudy (anonymous):

Refer to the form y=mx+b (m is your slope, and b is your y-value).

OpenStudy (anonymous):

I am very new into math ,tell me the steps and help me out

OpenStudy (anonymous):

The slope is 4 And the y intercept is 11. This is the formula y=mx+b y=4x+11 is already written in that formula so You don't have to change it. Just always remember the M stands for the SLOPE, and the B stands for the Y-INTERCEPT. It's pretty simple as long as you just remember the formula. (y=mx+b)

OpenStudy (anonymous):

By the way I'm new to this too ! (: